1. How many ways can three items be selected from a group of six items? Use the letters A, B, C, D, E, and F to identify the items, and list each of the different combinations of three items.


2. An experiment with three outcomes has been repeated 50 times, and it was learned that E1 occurred 20 times, E2 occurred 13 times, and E3 occurred 17 times. Assign probabilities to the outcomes. What method did you use?


1
Expert's answer
2021-05-20T18:26:22-0400
  1. Three items be selected from a group of six items in: C(6;3) = 6!/(6 - 3)!*3! = 20.
  2. Probabilities to the outcomes are: P(E1) = 20/50 = 0.4, P(E2) = 13/50 = 0.26, P(E3) = 17/50 = 0.34.